OpenStudy (anonymous):

I really need help with Linear Equations. I dont understand how someone can turn y=1/2x-5 into coordinates to go on a plane... That question above is just an example. I need some help understanding how to do this please.

terenzreignz (terenzreignz):

I need to understand your plight first... you need points to plot on the plane? ^_^

OpenStudy (phi):

\[ y = \frac{1}{2} x -5 \]?

OpenStudy (anonymous):

Yes in my lesson it is somehow adding points to that on a table?

terenzreignz (terenzreignz):

Well, if you're just looking for points, then the first step here is to just pick a value for x. Any value should work, but for simplicity, I recommend you pick even numbered values...

OpenStudy (phi):

the equation tells you (if you know how to interpret it) how to find a y value for a given x value.... the idea that for every x there is a special y that goes with that x

OpenStudy (anonymous):

So it can just be random? How is that math???

terenzreignz (terenzreignz):

You're looking at it the wrong way :) There are infinitely many points that would satisfy your equation, you certainly can't name them all. However, you can determine SOME points by picking a few values for x and determining the EXACT values of y that go with each respective x-value :P

OpenStudy (phi):

to find the y that goes with x=0 (for example) replace x in the equation with the number 0 y = (1/2)*0 - 5 and simplify. 1/2 times 0 is 0, so this is y = 0-5 and 0-5 is -5 all of that says that the pair (0,-5) go together

OpenStudy (phi):

yes, they are showing you how to do it. Can you do each of the steps ?

OpenStudy (anonymous):

Every thing but putting it into a table.

terenzreignz (terenzreignz):

Well, the points under the x-column were rather random too, chosen by whomever wrote that table... you can make your own table :)

OpenStudy (anonymous):

So i just pick a random x or a random y and work out the problem?

OpenStudy (phi):

the table is x and y in the x column you put x in the y column you figure out the y using that x

terenzreignz (terenzreignz):

You can pick a random x OR a random y-value, but not both. It's highly recommended that you pick x at random, as picking a random y-value would involve extra work at solving for x...(not by much though)

OpenStudy (phi):

you would pick an x (because if you pick the y, it will be hard to figure out the x unless you do algebra)

terenzreignz (terenzreignz):

lol, and as you can see, under the x-column, only even-number values were picked for x :P

OpenStudy (anonymous):

Ok i understand now. So if i was to do y=2x-10 I would pick my x y=2 * 5 - 10 and solve for y?

terenzreignz (terenzreignz):

Yup :)

OpenStudy (anonymous):

Ok TY guys so much i had no clue what i was doing XD!!!

terenzreignz (terenzreignz):

And you'll find that no matter what values you pick for x, they will all lie on the same line if you graph ^_^

OpenStudy (anonymous):

Oh Thank you i forgot about khanacademy